OpenStudy (anonymous):

a gas sample is heated from -20 degrees C to 57 degrees C and the volume is increased from 2 L to 4.5 L. If the initial pressure is 0.105 atm what is the final pressure

OpenStudy (aaronq):

you wanna use the ideal gas equation, PV=nRT, and discard anything that stays constant. so if you group everything to one side you end up with \(\dfrac{PV}{nRT}=0\) n and R are constant. so the equation you would use is: \(\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}\)

OpenStudy (anonymous):

Thank you!!

OpenStudy (aaronq):

no problem !